6. Mercury (I) nitrate should have the formula HgNO3, but spectroscopic and freezing point depression studies indicate (and i can you help with 6
Answer #1

Vant Hoff factor is the ratio between observed colligate property to the theoretically calculated colligative property.

Or it is ratio between observed number of particles to the theoretically calculated number of particles.

Vant Hoff factor is denoted by i

i = observed number of particles/Calculated number of particles

For HgNO​​​3 ,

Observed number of particles = 2 ( Hg​​​​​2+ and NO​​​​​3​​​​- )

Theoretically calculated number of particles = 1

Therefore i = 2/1 = 2

For Hg​​​​​2​​​​(NO​​3 )2 ,

Observed number of particles= 3 ( One Hg​​​​​2​​​​2+ and 2 NO​​​​​3​​​​- ions)

Theoretically calculated number of particles = 1

Vant Hoff factor for Hg​​​​​2​​​​(NO​​3 )2 is 3/1 = 3

i = 3/1 = 3